Find the sums of the following fractions.

1. 1/4 + 1/8

2. 1/2 + 3/10

3. 5/12 + 1/4

4. 2/5 + 4/15

5. 2/3 + 7/15

6. 2/3 + 7/15

7. 8/21 + 3/7


1
Expert's answer
2021-06-16T03:54:29-0400

1. 1/4 + 1/8

"\frac{1}{4}+ \frac{1}{8}= \frac{2+1}{8}= \frac{3}{8}"

2. 1/2 + 3/10

"\frac{1}{2}+ \frac{3}{10}= \frac{5+3}{10}= \frac{8}{10}"

3. 5/12 + 1/4

"\frac{5}{12}+ \frac{1}{4}= \frac{5+3}{12}=\frac{8}{12}=\frac{2}{3}"

4. 2/5 + 4/15

"\frac{2}{5}+ \frac{4}{15}= \frac{6+4}{15}=\frac{10}{15}= \frac{2}{3}"

5. 2/3 + 7/15

"\frac{2}{3}+ \frac{7}{15}= \frac{10+7}{15}= \frac{17}{15}"

6. 2/3 + 7/15

"\frac{2}{3}+ \frac{7}{15}= \frac{10+7}{15}= \frac{17}{15}"

7. 8/21 + 3/7

"\frac{8}{21}+ \frac{3}{7} = \frac{8+9}{21}=\frac{17}{21}"